# Scaling Laws for Economic Productivity: Experimental Evidence in LLM-Assisted Translation

### v1

@@ Abstract
This paper derives ‘scaling laws’ - empirical relationships between the amount of training compute used for a Large Language Model (LLM) and their performance- for economic outcomes. In a preregistered online experiment, 300 professional translators completed 1800 tasks with access to one of thirteen LLMs with differing model training compute sizes (or a control). Our results show that model scaling substantially raises productivity: for every 10x increase in model compute, translators completed tasks 12.3% quicker, received 0.18 s.d. higher grades, and earned 16.1% more per minute (including bonus payments). Further, the gains from model scaling are much higher for lower-skilled workers who gain a 4x larger improvement in task completion speed. These results imply further frontier model scaling- which is currently estimated at 4x increase per year- may have significant economic implications.
footnotetext: We gratefully acknowledge financial support from Open Philanthropy to conduct this research. The research described in this article was approved by the Yale Human Research Protection Program. We thank Charlie Harrison for his invaluable research assistance on this experiment.
@@ 1 Introduction
The amount of training compute used by frontier large language models (LLMs) increased by  5000x between the release of GPT-2 in 2019 and GPT-4 in 2023 and estimates from Epoch AI suggest a similar increase over the next six years. How does this massive increase in model training compute map onto performance? The empirical machine learning literature has derived remarkably consistent ‘scaling laws’ suggesting a strong relationship between a model’s training compute and model perplexity, a measure of model loss, across more than seven orders of magnitude. But there is so far a very limited understanding of how this reduction in perplexity affects key economic and social outcomes.
This paper aims to offer the first experimental evidence on this question by conducting a randomized controlled trial (RCT) involving 300 professional translators conducting 1800 tasks of varying complexities. The participants were randomly assigned to either treatment groups where they could utilize one of thirteen LLMs of differing model training compute to help them complete their task or to a control group where they completed tasks without any AI assistance. Participants face high-powered incentives with significant bonus payments for high-quality tasks as evaluated by three experienced professionals in the field. The key outcome variables, therefore, were how translator’s time taken, quality of tasks completed, and earnings per minute (inclusive of bonuses) varied by model training compute.
@@
The rest of the paper is organized as follows. Below, connections to four existing academic literatures are drawn. In Section Two, an overview of the experimental design is offered alongside some summary statistics on how the translator’s viewed the experimental design. This includes the degree to which they considered the tasks as similar to ones they experienced as professionals and their previous familiarity with AI tools. In Section Three, all the results are presented. Finally, in Section Four the paper concludes with discussion on the results.
@@
Thirdly, this paper also aims to contribute to a literature in labor economics aiming to understand the relationship between technological growth and skill-biased wage premia. Whether a given technology is skill-biased or not can have significant implications on changes in wage inequality as depicted by research by Goldin (2007). Preliminary experimental evidence from research in professional tasks as well as in the legal field suggest that using an LLM offers the greatest productivity enhancement to those who are lowest-skilled. This study aims to extend this literature by considering a continuous measure of automation to better understand how the skill-based wage premia changes according to increases in LLM capabilities.
Finally, this paper adds to the literature on the effect of machine translation on a translator’s productivity and quality. This literature has existed since 2013 (Kalchbrenner and Blunsom, 2013) and has seen a variety of machine tools used based on varying different architectures from neural machine translation (NMT) to the more recent transformer architecture. A good early literature review is provided by Sanchez and Torron (2017). This paper contributes to this literature by analyzing the effects of thirteen different LLMs of different compute sizes on translator’s productivity and quality.
@@ 2 Experimental Design
Over 70% of eligible participants had at least three years of professional experience as a translator and just over a third had at least five years of experience as a professional. In self-reported analysis, translator’s judged their translation abilities between English and the target language at 4.52/5 and their professional fluency in the desired language at 4.85/5. Further, they reported a familiarity of AI tools for translation of 4.15/5 and self-evaluated their abilities in using AI tools for translation at 4.18/5.

### v2

# Ali Merali

@@ Abstract
This paper derives ”scaling laws”—empirical relationships between the training compute of Large Language Models (LLMs) and their performance—for economic outcomes. In a preregistered online experiment, 300 professional translators completed 1,800 tasks using one of 13 LLMs (or a control). A tenfold increase in model compute improved task completion speed by 12.3%, grades by 0.18 standard deviations, and earnings per minute by 16.1%. Gains were four times larger for lower-skilled workers. These findings suggest continued model scaling could boost U.S. productivity by at least 6.9% over the next decade.
footnotetext: Yale University, Department of Economics: ali.merali@yale.edu
@@ 1 Introduction
The amount of training compute used by frontier large language models (LLMs) increased by roughly 5000x between the release of GPT-2 in 2019 and GPT-4 in 2023. Estimates from Epoch AI, an AI research institute, suggest a similar increase over the next six years. How does this massive increase in model training compute, however, map onto performance? The empirical machine learning literature has derived remarkably consistent ‘scaling laws’ suggesting a strong relationship between a model’s training compute and model perplexity, a measure of model loss, across more than seven orders of magnitude. But there is so far a very limited understanding of how this reduction in perplexity affects key economic and social outcomes.
This paper aims to offer the first experimental evidence on this question. It conducts a randomized controlled trial (RCT) involving 300 professional translators conducting 1800 tasks of varying complexities. The participants were randomly assigned to either treatment groups where they could utilize one of thirteen LLMs of differing model training compute to help them complete their task or to a control group where they completed tasks without any AI assistance. Participants faced high-powered incentives with significant bonus payments for high-quality tasks as evaluated by three experienced professionals in the field. The key outcome variables, therefore, were how translator’s time taken, quality of tasks completed, and earnings per minute (inclusive of bonuses) varied by model training compute.
@@
Utilizing the experimental findings on economic scaling laws, I estimate the aggregate productivity gains from AI on the US economy over the next decade. This calculation mirrors that of Acemoglu (2024) and utilizes the same framework based on Hulten’s theorem- the only changes are to allow for the possibility of model improvement as the scaling hypothesis suggests, and to take into account recent data on inference costs. These changes leads to a significantly higher productivity estimate of around a magnitude higher than in Acemoglu (2024) with an estimate 6.9% productivity growth.
The rest of the paper is organized as follows. Below, connections to four existing academic literatures are drawn. In Section Two, an overview of the experimental design is offered alongside some summary statistics on how the translator’s viewed the experimental design. This includes the degree to which they considered the tasks as similar to ones they experienced as professionals and their previous familiarity with AI tools. In Section Three, all the results are presented. In Section Four, a basic quantitative estimate is made for the aggregate productivity gains from AI over the next decade. Finally, in Section Five the paper concludes with discussion on the results.
@@
Thirdly, this paper also aims to contribute to a literature in labor economics aiming to understand the relationship between technological growth and skill-biased wage premia. Whether a given technology is skill-biased or not can have significant implications on changes in wage inequality as depicted by research by Goldin (2007). Preliminary experimental evidence from research in professional tasks as well as in the legal field suggests that using an LLM offers the greatest productivity enhancement to those who are lowest-skilled. This study aims to extend this literature by considering a continuous measure of automation to better understand how the skill-based wage premia changes according to increases in LLM capabilities.
Finally, this paper adds to the literature on the effect of machine translation on a translator’s productivity and quality. This literature has existed since 2013 (Kalchbrenner and Blunsom, 2013) and has seen a variety of machine tools used based on varying different architectures from neural machine translation (NMT) to the more recent transformer architecture. A good early literature review is provided by Sanchez and Torron (2017). This paper contributes to this literature by analyzing the effects of 13 different LLMs of different compute sizes on translator productivity and quality.
@@ 2 Experimental Design
Over 70% of eligible participants had at least three years of professional experience as a translator and just over a third had at least five years of experience as a professional. In self-reported analysis, translators judged their translation abilities between English and the target language at 4.52/5 and their professional fluency in the desired language at 4.85/5. Further, they reported a familiarity of AI tools for translation of 4.15/5 and self-evaluated their abilities in using AI tools for translation at 4.18/5.
@@ 3 Results
4 Aggregate Productivity Gains
In this section, we utilize the above results on the productivity gains from model scaling to estimate aggregate productivity gains from AI over the next ten years. In doing so, we leverage the framework from Acemoglu (2024). Here a version of Hulten’s theorem is used to estimate aggregate productivity gains based on the fractions of tasks in the economy that are affected by AI and the average task-level cost savings.
The estimates from Acemoglu (2024) therefore are based on the multiplication of four parameters. Firstly, an estimate of the share of tasks exposed to AI is used from Eloundou et al (2023) which estimates it at 19.9%. Secondly, as some tasks involve a combination of labor and capital the (AI-exposure adjusted) labor share of 0.57 is used. Neither of these parameters are changed in the analysis of this section.
However, the third parameter estimating the labor cost savings from AI is changed significantly. The estimate from Acemoglu (2024) come from two studies; Brynjolfsson et al (2023) which looked at a real-world case study involving call center operatives and Noy and Zhang (2024) who used an online experiment for a range of white-collar tasks. The estimate comes from an average of the two productivity effect sizes which is 27%. Note that these estimates do not take into account the possibility that the productivity effects may grow over time as models scale in the amount of training compute used and model quality improves. Further, they combine estimates from two different generations of models; the lower estimate of 14% from Brynjolfsson et al (2023) comes from the usage of a GPT-based model rolled out in 2020-2021 whereas the 40% estimate in Noy and Zhang (2023) comes from the larger and higher quality ChatGPT model (both models have already been superseded by much larger and higher-performing models by OpenAI, Anthropic, Meta, Alphabet, and others in 2024 alone).
In particular, the 14% estimate comes from a model estimated to be trained on the order of $10^{23}$ training FLOPs of model compute whilst the 40% estimate comes from a model estimated to be trained on the order of $10^{24}$ training FLOPs. This compares to models trained on the order of $10^{25}$ training FLOPs in this study. It is likely that if either of these studies was completed on the best available models today the productivity effects would be significantly higher - how can we estimate just how large we expect these productivity estimates to be?
Here we use evidence from the scaling results from this experiment. Note we do not assume that the scaling laws for different tasks all follow those of the scaling laws for translation- other tasks may see differential productivity gains. The weaker assumption that is made here, however, is that the ratio of the productivity gains made from going from no AI assistance to assistance from an LLM with training compute x to the productivity gains from going from assistance with an LLM with training compute x to an LLM with a larger training compute y for any given task is the same as it is for translation.
Note that Eloundou et al (2023) estimate translators to be the occupation with either the highest or close to the highest (depending on the model they use) AI-exposed occupation. As such, we would expect the ratio of gains from AI-assistance to be disproportionately larger from smaller models to larger models for translators than for other tasks. This assumption, therefore, likely yields an underestimate for the scaling effects of all other tasks.
Nonetheless, evidence from Table 2 suggests that for translation 54.9% of the gains from using the largest AI model in this study, Claude-3.5 Sonnet, would be received by a model of the size used in Brynjolfsson et al (2023) and 72.8% of the gains from the model of the size used in Noy and Zhang (2023). This therefore estimates that if these studies were replicated with the largest currently available models their productivity effects would be 25.5% and 54.9% respectively for an average of 40.2%.
Of course, the current largest models are extremely unlikely to be the largest models available in ten years time. Conservatively, we use Epoch AI’s estimates of models with the order of $10^{30}$ training FLOPs being commercially available by 2030 and allow the remaining time for their widespread adoption. Again utilizing evidence from Table 2, and assuming the ratio of progress between the jump from no models to smaller models and the jump from smaller models to larger models mirrors the estimates for translation, we estimate that replications of Brynjolfsson et al (2023) with the best models available in 2030 would have a 39.1% productivity boost whilst similar replications of Noy and Zhang (2023) would see a 84.4% boost. This yields an average of a 61.2% productivity effect in ten years time when we account for (both past and) future model scaling, compared to the original 27% estimate.
Finally, the remaining parameter in Acemoglu’s estimate is whether tasks that can be automated can be done so profitably. Here, he uses an estimate from Svanberg et al (2024) suggesting 23% of automatable tasks would be economically feasible to do so within then years. However, this estimate is for computer vision tasks which can involve high fixed costs such as camera equipment. However, the cost can be substantially smaller for tasks that can be automated through LLMs. Indeed the inference costs for all 300 translators completing nearly 1,400 tasks with AI assistance combined came to significantly less than a dollar.  [^1] This ignores the time cost of a professional interacting with the LLM interface, by far a more significant cost than inference costs, however this is already accounted for in the time-based productivity estimates For this domain therefore, the profitable automation rate would be roughly 100%.
More generally, the inference cost of one of the leading models as of the time of writing is \$1 for $100,000$ syllables of writing. As such, if any task can be automated via an LLM it is almost certain it will be profitable to do so based on inference costs. Further, a report from Andreesen Horowitz, a venture capitalist firm, shows that the inference cost of an LLM achieving a certain fixed benchmark score has recently been falling by 10x per year. If it wouldn’t currently be profitable to use an LLM to aid for a task, therefore, it is likely to be so in the near-term future even if such a rapid rate of cost decreases doesn’t continue. This means that for the purpose of this calculation we assume that if LLMs can increase productivity on a task it will always be profitable.
This means that the total productivity gains from AI over the next ten years is estimated through the multiplication of three parameters. The share of tasks that can be automated (19.9%) multiplied by the average productivity effect using scaling laws (61.2%) multiplied by the labor share of costs (57%). This yields an estimate of 6.95% productivity growth over the next decade from LLMs, a significant amount of growth.
It should be noted that this calculation makes several restrictions. For instance, it is assumed that the task structure of the economy doesn’t change at all despite significant technological changes. Further, and of particular significance, it also assumes no change to the rate of technological progress itself. There is already a wealth of evidence, however, that AI can increase the pace of scientific discovery itself for instance in predicting protein structures (Jumper et al 2021) or discovering new materials technology (Toner-Rodgers, 2024). This can lead to much higher estimates such as in Korinek and Suh (2024), therefore there is a sense in which these estimates may be considered a lower bound.